OpenStudy (anonymous):

What is the value of the expression 2 exponent 4 + 3 exponent 2?

OpenStudy (photon336):

\[2^{4}+3^{2} = ? \]

OpenStudy (anonymous):

yeah that really helps #sarcastic

OpenStudy (rawfeeds-):

you want him or her to do the work for you do it yourself if you dont want help

OpenStudy (photon336):

If I gave you the direct answer that wouldn't help either now wouldn't it?

OpenStudy (igreen):

\(\sf 2^4 \rightarrow 2 \times 2 \times 2 \times 2\) \(\sf 3^2 \rightarrow 3 \times 3\) Can you multiply those two and add them together?

OpenStudy (anonymous):

yeah thankyou umm the first one is 16 and the other is 9 so if u add them together it would be 25?

OpenStudy (igreen):

Yes, that's correct. Just remember what exponents mean.

OpenStudy (anonymous):

ok like multiply the same number how many times the exponent is